after tax cost of debt

Burnwood Tech plans to issue some $60 par preferred stock with a 6% dividend. A similar stock is selling on the market for $70. Burn wood must pay flotation costs of5% of the issue price. What is the cost of the prferred stock?
Answer #1
Par Value of Preferred Stock = $60
Dividend percentage = 6%
Market Value of the stock = $70
Flotation costs = 5% of the issue price
Cost of Preferred Stock =?


Annual Dividend Payment on Preferred Stock = [6% * $60]
Annual Dividend Payment on Preferred Stock = $3.60

Cost of Preferred Stock = [Preferred Stock dividend / Market Price of
Preferred Stock]
[OR]

Cost of Preferred Stock = [Preferred Stock dividend / Market Price of
Preferred Stock (1-Flotation cost)]

Cost of Preferred Stock = [($60 * 6%) / $70 (1-0.05)]
Cost of Preferred Stock = [$3.60 / $70 (0.95)]
Cost of Preferred Stock = [$3.60 / $66.50]
Cost of Preferred Stock = 0.054135

Cost of Preferred Stock = 5.41%
answered by: Richie

> There needs to be a parenthesis in the denominator. The correct formula should show:

Cost of Preferred Stock =(60*0.06)/(70*(1-0.05))
Cost of Preferred Stock =3.60/(70(0.95))
Cost of Preferred Stock = 3.60/66.50
Cost of Preferred Stock = 0.054135

Frank Abagnale Mon, Nov 22, 2021 9:05 PM